OpenStudy (anonymous):

Simplify completely quantity 6 x squared minus 54 plus 84 over quantity 8 x squared minus 40 x plus 48 divided by quantity x squared plus x minus 56 over quantity 2 x squared plus 12 x minus 32

OpenStudy (anonymous):

a. 2 open parentheses x minus 3 close parentheses over 3 open parentheses x minus 2 close parentheses b. 2 open parentheses x minus 2 close parentheses over 3 open parentheses x minus 3 close parentheses c. 3 open parentheses x minus 3 close parentheses over 2 open parentheses x minus 2 close parentheses d. 3 open parentheses x minus 2 close parentheses over 2 open parentheses x minus 3 close parentheses

OpenStudy (anonymous):

You should be able to answer this question the same way as the last question.

jhonyy9 (jhonyy9):

try factorizing these quadratic equations and than will see what you will can simplify

jhonyy9 (jhonyy9):

6x^2 -54x +84 = 6(x^2 -9x +14) = 6(x -7)(x-2) 8x^2 -40x +48 = 8(x^2 -5x +6) = 8(x-3)(x-2) 2x^2 +12x -32= 2(x^2 +6x -16) = 2(x+8)(x-2) x^2 +x -56 = (x+8)(x-7) 6x^2 -54x +84 --------------- 8x^2 -40x +48 6x^2 -54x +84 2x^2 +12x -32 ----------------- = --------------- * ---------------- = x^2 +x -56 8x^2 -40x +48 x^2 +x -56 --------------- 2x^2 +12x -32 6(x -7)(x-2) 2(x+8)(x-2) 3(x-7) 2(x-2) 3*2(x-2) = ------------ * ------------ = ------- * ------- = --------- = 8(x-3)(x-2) (x+8)(x-7) 4(x-3) x -7 4(x-3) 3(x-2) = -------- 2(x-3) hope so much that is understandably sure right easy good luck bye